What is the resultant and the angle of the coplanar forces which is 100 N at 30°, 141.4 N at 45 ° and 100 N at 240 °? Show your solution and draw position of the resultant in the x-y plane. Final answer must be in 3 significant figures.
1
Expert's answer
2020-11-02T09:23:08-0500

Find the components:


"F_{1x}=F_1\text{cos}30°=86.6\text{ N},\\
F_{1y}=F_1\text{sin}30°=50\text{ N}.\\
F_{2x}=F_2\text{cos}45°=99.98\text{ N},\\
F_{2y}=F_2\text{sin}45°=99.98\text{ N}.\\
F_{3x}=F_3\text{cos}240°=-50\text{ N},\\
F_{3y}=F_3\text{sin}240°=-86.6\text{ N}.\\"

The resultant is


"F_x=F_{1x}+F_{2x}+F_{3x}=137\text{ N},\\
F_y=F_{1y}+F_{2y}+F_{3y}=63.4\text{ N},\\
F=\sqrt{F_x^2+F_y^2}=151\text{ N},\\\space\\
\theta=\text{arctan}\frac{F_y}{F_x}=24.9°"

above x-axis.